6. a) Given any real symmetric square matrix A of order n, show that the function RnxRn "\to" defined by I(x, y) =xTAy is a symmetric bilinear function

Give examples of matrices A for which the corresponding I is not an inner product

State the extra properties on A which would make the corresponding I an inner product


b) Show that R[x] deg ≤ 2 x R[x] deg ≤ 2 defined by I (p(x), q(x)) =p(0)q(0)+p(1/2)q(1/2)+p(1)q(1) is an inner product on R[x] deg ≤ 2

Use Gram-Schmidt procedure to obtain a orthogonal basis of R[x] deg ≤ 2(with respect to the inner product defined above) which contains the polynomial x


1
Expert's answer
2021-11-15T05:09:39-0500

The function defined by "\Iota(x,y) = x^TAy" is a symmetric bilinear function

1) it satisfies the axiom <u,v> = <v,u> in such:


⟨x,y⟩=x"^T" Ay=x⋅(Ay)


Since A is symmetric


⟨x,y⟩=x⋅(Ay)=(Ay)⋅x

=(Ay)"^T"x=y"^T" A"^T" x=y"^T" Ax=⟨y,x⟩.


(ii) For any vector x,y,z and any real number r, we have:


"\langle rx,y\rangle=(rx)^TAy=rx^TAy=r\langle x,y\rangle"

and

"\langle\>x+y,z\rangle=(x+y)^TAz=(x^T+y^T)Az"


"=x^TAz+y^TAz=\langle\>x,y\rangle=\langle\>y,z\rangle"


Therefore the linearity in the first argument is satisfied


It satisfies the rule of positive- Definiteness in such


If "x\ne0" then we have

"\langle\>x,x\rangle=x^TAx>0"


Examples of matrices A for which the corresponding "\Iota" is not an inner product

"\to \>" A matrix of order m"\u00d7"n

"\to" A matrix A "\ne\>A^T"

"\to" A matrix A whose eigenvalues "\lambda\> \leq0"

"\to" A matrix A whose determinant det "A=0"


Properties of A which would make the corresponding "\Iota" on inner product

Matrix A should be a real symmetric "n\u00d7n" matrix

Matrix A must be positive definite in that

"x^TAx>0"

Matrix A has real eigenvalues

The eigenvectors of matrix A corresponding to the eigenvalues are orthogonal

Matrix A is diagonalizable


Part 6(B) sections


"\Iota" is inner product if it satisfies conjugate symmetry, linearity and positive Definiteness axioms

(I) "<q(x),p(x)>=q(0)p(0)+q(\frac{1}{2})q(\frac{1}{2})+q(1)p(1)"


"=p(0)q(0)+p(\frac{1}{2})q(\frac{1}{2})+p(1)q(1)"


"\therefore<q(x),p(x)>=<p(x),q(x)>"


"\Iota" is symmetric


(ii) "c<p,q>=c\>[{p(0)q(0)+p(\frac{1}{2}q(\frac{1}{2}+p(1)q(1)}]"


"=cp(0)q(0)+cp(\frac{1}{2})q(\frac{1}{2})+cp(1)q(1)"


"=<cpf,q>"

"\Iota" satisfies linearity


(iii) "<p,p>=p(0)p(0)+p(\frac{1}{2})p(\frac{1}{2})+p(1)p(1)>0"


If "p(x)" is not zero


Part 6(B) section 2


The standard polynomial basis for "R[x]deg\>\le2" is defined "[1,x,x^2]"

Let orthogonal basis of "\Iota" be "(b_1,b_2,b_3)"

"b_1=1"

"b_2=x-proj_1x"


"\therefore b_2=x-\frac{<1,x>}{<1,1>}1"


"<1,x>=(1)(0)+(1)(\frac{1}{2})+(1)(1)=\frac{3}{2}"


"<1,1>=(1)(1)+(1)(1)+(1)(1)=3"


"b_2=x-\frac{1}{2}"


"b_3=x^2-\frac{<x^2,1>}{<1,1>}1-" "\frac{<x^2,x-\frac{1}{2}>}{<x-\frac{1}{2},x-\frac{1}{2}>}(x-\frac{1}{2})"



"<x^2,1>=(1)(0)+(1)(\frac{1}{2})^2+(1)(1)^2=\frac{5}{4}"


"<x-\frac{1}{2},x-\frac{1}{2}>=(\frac{-1}{2})(\frac{-1}{2})+(0)(0)+(\frac{1}{2})(\frac{1}{2})=\frac{1}{2}"



"<x^2,x-\frac{1}{2}>=(0)(\frac{-1}{2})+(^\frac{1}{2})^2(0)+(1)^2(\frac{1}{2})"


"=\frac{1}{2}"


"b_3=x^2-\frac{5}{12}-1(x-\frac{1}{2})"


"=x^2-x+\frac{1}{12}"